OpenStudy (anonymous):

Please help me out. :) Find the maximum and minimum then graph. y = 3 - 3x^2 - x^3

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

No. That's my prof's idea.

OpenStudy (loser66):

take derivate of y, let it =0 solve for x. Show me your work

OpenStudy (anonymous):

dy/dx = -6x - 3x^2 0 = -6x - 3x^2 0 = -3x (2+x) 0 = -3x and 0 = 2+x x = 0 and x = -2

OpenStudy (loser66):

so, what are f(0) and f(-2)?

OpenStudy (anonymous):

f(0) = 3 f(-2) = -1

OpenStudy (loser66):

do you know how to make the table? like